# Effects of lower floating-point precision on scale-resolving numerical simulations of turbulence

### v1

@@ 5 Transition to turbulence
We consider so-called K-type transition where a laminar baseflow as described by Schlatter et al.  [48, 49] is perturbed by one 2D and two oblique 3D Tollmien–Schlichting (TS) waves with amplitudes of $3\%$ and $0.1\%$ respectively (based on the laminar centerline velocity), all of which are individually stable. K-type (Klebanoff) transition refers to the instability being of fundamental type, i.e. the streamwise wavenumber of the primary and secondary instability are the same. The eigenvectors for the initial condition were computed using a Jupyter notebook [50] and superposed on top of the parabolic laminar Poiseuille flow. The Reynolds number is $Re_{b}=3333$ based on the constant bulk velocity, which corresponds to $Re_{cl}=5000$ based on the centerline velocity $U_{cl}$ of the initial parabolic velocity profile and the channel half-height $h$. The domain size is $5.61\delta\times 2.99\delta\times 2\delta$, adjusted to fit the chosen TS waves with $\alpha_{0}=1.12$ and $\beta_{0}=2.1$ as the streamwise and spanwise fundamental wavenumbers. For the compressible codes, instead of fixed mass flux, a constant pressure gradient forcing is applied in the streamwise direction to drive the flow. This would lead to a lower turbulent Reynolds number, but the initial growth of perturbations are only marginally affected. Thus, in all cases, a matching bulk Reynolds number of $Re_{b}=3333$ is maintained before turbulent breakdown.
@@
We first focus on the expected behavior during transition, as illustrated in Figure 6. Panel a) shows the evolution of the two-dimensional (spanwise) modes $|\hat{u}(\alpha,\beta=0)|$, for integer $\alpha=0,1,\ldots$. It is always the maximum absolute value of the mode over the channel shown. It is clear that the mean-flow modes $\alpha=0$ and $\beta=0$ is only changing at $t>150h/U_{cl}$ corresponding to the establishment of the turbulent profile with a lower centerline component. The only other non-zero mode at $t=0$ is the 2D TS wave with $3\%$ energy. However, due to the nonlinearity of the flow and the triadic interactions, the flow quickly establishes a saturated 2D TS wave, with higher and higher 2D modes being energized, with a weak temporal decay.
@@
Secondary instability, initiated by the $\beta=1$ modes, leads to a quick increase in the energy in all modes ($t>120h/U_{cl}$), the formation of characterisitic hairpin vortices ($t=136h/U_{cl}$, see Figure 7 )and subsequent breakdown to turbulence ($t>175h/U_{cl}$). The double-precision arithmetic allows us to resolve numerically all modes down to machine precision ($10^{-15}$) for Simson, but saturates at around $10^{-9}$ for the other solvers. Reducing to single precision increases the ambient noise level to about $10^{-8}$ for Simson, and around one order of magnitude higher for the other codes. Interestingly, there seems to be no interaction between these modes that would lead to a premature growth in the physically relevant modes. In contrast, similar studies using low-resolution simulations found a clear change of energy distribution and subsequent growth, which can be contained only using appropriate subgrid-scale models [48, 49]. It is noteworthy to highlight that not only the Fourier amplitudes and integral quantities are seemingly not affected by the lower precision, but also the actual vortical flow structures, as shown in Fig. 7. As opposed to lower-resolution simulations (as cited previously), turbulence is not appearing prematurely or disturbing the flow. The sharp gradients around the hairpin heads are well resolved without artifacts.
@@
From Figure 8 we can conclude that the evolution of the individual modes, but also integral quantities such as the global friction or centerline velocities, is not dependent on the precision. For the rounding of the state and convective terms in the different solvers, we also observe that FP32 performs remarkably well, but when representing the state at lower-precision, the simulation becomes prone to stagnation (horizontal lines) or an immediate transition (State FP16). However, although the transitional case is sensitive, the amplitude of the initial conditions is still on the order of 0.10.10.10.11111%, and there will be a precision-dependent limit on the smallest disturbance amplitude the simulations would be able to capture. In addition, the geometry is a Cartesian channel, which motivates the study of a deformed geometry, such as the separating flow around a cylinder.
@@
Figure 6: Evolution of Fourier modes during K-type transition, obtained with Simson. The solid lines are FP64 (double precision), and the black dotted lines FP32 (single precision). a) shows the two-dimensional Fourier modes $|\hat{u}(\alpha,\beta=0)|$, starting with $\alpha=0$ on the top. (b) Three-dimensional Fourier modes $|\hat{u}(\alpha=1,\beta)|$ starting with $\beta=0$ on the top. The wall-normal maximum is shown in both cases.
@@
Figure 8: Evolution in time of amplitude of 2D modes $|\hat{u}(\alpha,\beta=0)|$ for transitional case. All simulations carried out in Neko and SSDC, except the reference case in Simson, shown in gray. Results from full FP32 and FP64 (a), State rounding (b), and rounding of the convective term (c) are shown.
Figure 9: Profiles for the cylinder at $Re_{D}=3900$ in Neko and SSDC with rounding the convective term, the state, and running the entire solver in single and double precision. The $C_{p}$ profile in the center of wake (a) and wake profile at four different locations in the wake $(0.58,1.06,1.54,2.02)$ (b), the blue-shaded interval is between time averages of two low-frequency modes as described by [51].
@@
Setup Avg. time ($D/U_{\infty}$) $f_{vs}$ $\phi_{s}$ $L_{r}$ $\bar{C}_{d}$ $-\bar{C}_{Pb}$
[52] 35 0.21 88 1.35 1.04 0.94
@@ 6 Separated flow — Cylinder at $Re_{D}=3900$
This section considers the flow around an infinite circular cylinder at $Re_{D}={U_{\infty}D}/{\nu}=3900$, where $D$ is the cylinder diameter and $U_{\infty}$ the free-stream velocity. We perform LES in PadeLibs, SSDC and Neko with approximately 512 grid points along the cylinder boundary, and a spanwise length of $2\pi D$ with $128$ grid points. There is extensive literature on this case, showing a significant spread in the simulation results [51]. The results are illustrated in Figure 9, which shows the velocity profiles in the wake and pressure distribution on the cylinder surface. Table 4 compares these results with the original LES by [52] and highlights wall quantities such as the drag coefficient and separation angle, as well as the length of the recirculation zone. Overall, the differences among the setups and solvers are comparable to the spread in the reference data. As such, for the simulations that do not diverge, this case indicates that other sources of uncertainty are more significant than the numerical precision when it comes to LES of separating flows such as the size of the domain and, in particular, averaging times.
Isolating the impact of lower precision might become clearer with longer averaging times, but due to the discrepancies among multiple reference data it is not clear whether the impact of precision can be isolated. A case with a stronger consensus among the reference data and where shorter averaging times are necessary would likely be better suited to evaluate only the impact of numerical precision.
For a more nuanced comparison, we employ the Wasserstein distance of the probability density functions of the drag coefficient $C_{D}$. The Wasserstein distance is a metric to compare the similarity between PDFs and has previously been employed for a similar purpose (impact of rounding and floating-point precision) in [6] for climate models where the similarity of the statistical description of the system is under consideration.
By comparing the Wasserstein distance between the drag coefficient $C_{D}$ for the different cases, the impact of averaging times is further amplified. We show the distance between the probability distributions of $C_{D}$ among the different runs in Figure 10. In this plot, the variability for the runs which were only carried out for 100 time units are clearly evident. For Neko, it is only for the longer averaging times that FP32 and FP64 consistently have a smaller Wasserstein distance than the poorer averaged results. While the convective FP32 might appear to differ, by comparing the longer FP64 run with exactly the same setup but averaged for a shorter time, we also see a large difference that is aligned with the low-frequency oscillations that are also described in [51].
@@ 7 Compressible flow around a wing section
[…]
This flow regime is characterized by a laminar separation bubble with transition and turbulent reattachment, as illustrated by plotting iso-contours of the second invariant of the velocity gradient tensor in Figure 11. The flow is initialized with a preliminary two-dimensional solution, the simulation is run in FP64 for $15$ convective time units ($C/U_{\infty}$), after which the statistics are computed for an extra $12$ convective time units for both considered precisions.
Table 5 compares the averaged integral loads and the separation bubble extension between the two SSDC runs and the reference data. Figure 12 shows the mean pressure coefficient and skin friction coefficient on the airfoil surface. Overall, single precision computations achieve similar results compared to double precision, and both compare well to the reference. Minor differences with respect to the reference are visible in the skin friction coefficient plot, close to the leading edge on the pressure side. These can likely be attributed to a slightly different discretization of the laminar boundary layer, which is very thin in this region. Figure 13 shows mean velocity profiles and the separation bubble along the suction side for both precisions. Both computations achieve very similar results even for these quantities. The time dependence of the separation is assessed considering the probability density function of the skin friction coefficient along the suction side of the airfoil, Figure 14. In particular, the same procedure as in [53], Fig. 13c, has been applied for e.g. binning. Even for this sensitive quantity, both solutions demonstrate remarkably similar results.
@@
Setup $t_{avg}U_{\infty}/C$ $C_{L}$ $C_{D}$ $C_{D_{P}}$ $C_{D_{sf}}$ $\text{x}_{1}|_{C_{f}=0}$ $\text{x}_{2}|_{C_{f}=0}$
@@
Table 5: Scalar aerodynamic results associated with the airfoil at $Re_{C}=50000$ and $M=0.5$. Columns correspond to each setup name, associated time-averaging duration (in $C/U_{\infty}units$), the lift coefficient $C_{L}$, the drag coefficient $C_{D}$ (split into pressure and skin friction components), and the start and end points of the separation bubble, $\text{x}_{1,2}|_{C_{f}=0}$.
@@
Figure 14: Contour plot of the PDF of friction coefficient $C_{f}$ along airfoil chord. The PDF is scaled given a non-uniform bin size along $x$, more details can be found in [53] (Eq. 3.1, Fig. 13).
@@ 8 Practical experiences
This section summarizes our experiences in running our solvers natively in FP32. We comment on both performance gains (particularly on GPUs) and a range of issues that needed to be addressed for the solvers to run problem-free. While the identified problems differ among the solvers, collectively they serve as useful pointers to what should be treated with extra care when implementing an FP32-capable solver.
@@ 8.1 Performance improvements with FP32 over FP64
One of the most important practical outcomes from the simulation campaign was evaluating the performance impact of using FP32 instead of FP64 for the different test cases. Recall that F32 was the only precision for which the rounding was available directly on the hardware and not emulated, thus for the other precisions such measurements were not possible to obtain.
We gained most experience from our runs with Neko where a performance boost around 2×2$\times$2 × was expected as most kernels operate in the memory-bound domain. However, it was observed that the performance improvement changed depending on whether server-grade or consumer GPUs were used, as illustrated in Table 6.
In general, it was found that a larger problem size was needed to obtain 2×2$\times$2 × performance improvement on the server grade GPUs going from FP64 to FP32. For the cylinder case, for example, running on the AMD Instinct MI250X GPUs the performance was more or less exactly 2×2$\times$2 × faster with single precision. However, when a smaller turbulent channel flow case was considered this was no longer as evident as the problem size was insufficient to hide the latency of kernel launches and oversubscribe the available computational resources.
A comparison between the server-grade Nvidia A100 and the consumer-level Nvidia RTX 4080 revealed that, for single precision runs, the A100 achieved a performance increase of approximately 1.5×1.5$\times$1.5 × over its double precision counterpart in the smaller channel flow case. In contrast, the RTX 4080 demonstrated more than double the performance in single precision, and its FP32 performance matched that of the much more expensive A100 operating in FP64. This suggests that it is feasible to perform DNS on consumer-grade GPUs using FP32, achieving comparable performance to traditional FP64 simulations on high-end server GPUs. In this case, simulations executed in FP32 on the RTX 4080 produced results equivalent to those previously obtained with FP64 on the A100—within the same runtime.
@@
RTX4080 0.72 TB/s 0.76 TFlop/s 48.7 TFlop/s 22.5×2-2.5$\times$2 - 2.5 ×
A40 0.7 0.58 37.4 22.5×2-2.5$\times$2 - 2.5 ×
A100 1.56 9.7 19.4 1.52×1.5-2$\times$1.5 - 2 ×

### v2

@@ 5 Transition to turbulence
We consider the so-called K-type transition where a laminar baseflow, as described by Schlatter et al.  [51, 52], is perturbed by one 2D and two oblique 3D Tollmien–Schlichting (TS) waves with amplitudes of $3\%$ and $0.1\%$ respectively (based on the laminar centerline velocity), all of which are individually stable. K-type (Klebanoff) transition refers to the instability being of fundamental type, i.e. the streamwise wavenumbers of the primary and secondary instability are the same. The eigenvectors for the initial condition were computed using a Jupyter notebook [53] and superposed on top of the parabolic laminar Poiseuille flow. The Reynolds number is $Re_{b}=3333$ based on the constant bulk velocity, which corresponds to $Re_{cl}=5000$ based on the centerline velocity $U_{cl}$ of the initial parabolic velocity profile and the channel half-height $h$. The domain size is $5.61\delta\times 2.99\delta\times 2\delta$, adjusted to fit the chosen TS waves with $\alpha_{0}=1.12$ and $\beta_{0}=2.1$ as the streamwise and spanwise fundamental wavenumbers. For the compressible codes, instead of a fixed mass flux, a constant pressure gradient forcing is applied in the streamwise direction to drive the flow. This would lead to a lower turbulent Reynolds number, but the initial growth of perturbations is only marginally affected. Thus, in all cases, a matching bulk Reynolds number of $Re_{b}=3333$ is maintained prior to turbulent breakdown.
@@
We first focus on the expected behavior during transition, as illustrated in Figure 6. Panel a) shows the evolution of the two-dimensional (spanwise) modes $|\hat{u}(\alpha,\beta=0)|$, for integer $\alpha=0,1,\ldots$. It is always the maximum absolute value of the mode over the channel shown. The mean-flow modes $\alpha=0$ and $\beta=0$ are only changing at $t>150h/U_{cl}$, corresponding to the establishment of the turbulent profile with a lower centerline component. The only other non-zero mode at $t=0$ is the 2D TS wave, which has $3\%$ energy. However, due to the nonlinearity of the flow and the triadic interactions, the flow quickly establishes a saturated 2D TS wave, with higher and higher 2D modes being energized, with a weak temporal decay.
@@
Secondary instability, initiated by the $\beta=1$ modes, leads to a quick increase in the energy in all modes ($t>120h/U_{cl}$), the formation of characterisitic hairpin vortices ($t=136h/U_{cl}$, see Figure 7 )and subsequent breakdown to turbulence ($t>175h/U_{cl}$). The double-precision arithmetic allows us to resolve numerically all modes down to machine precision ($10^{-15}$) for Simson, but saturates at around $10^{-9}$ for the other solvers. Reducing to single precision increases the ambient noise level to about $10^{-8}$ for Simson, and around one order of magnitude higher for the other codes. Interestingly, there seems to be no interaction between these modes that would lead to a premature growth in the physically relevant modes. In contrast, similar studies using low-resolution simulations have found a clear change in energy distribution and subsequent growth, which can be contained only using appropriate subgrid-scale models [51, 52]. It is noteworthy to highlight that not only the Fourier amplitudes and integral quantities are seemingly not affected by the lower precision, but also the actual vortical flow structures, as shown in Fig. 7. In contrast to lower-resolution simulations (as cited previously), turbulence does not appear prematurely or disrupt the flow. The sharp gradients around the hairpin heads are well resolved without artifacts.
@@
From Figure 8, we can conclude that the evolution of individual modes, as well as integral quantities such as the global friction or centerline velocities, is not dependent on the precision. For the rounding of the state and convective terms in the different solvers, we also observe that FP32 performs remarkably well, but when representing the state at lower-precision, the simulation becomes prone to stagnation (horizontal lines) or an immediate transition (State FP16). However, although the transitional case is sensitive, the amplitude of the initial conditions is still on the order of 0.10.111%, and there will be a precision-dependent limit on the smallest disturbance amplitude the simulations can capture. In addition, the geometry is a Cartesian channel, which motivates the study of a deformed geometry, such as the separating flow around a cylinder.
@@
Figure 6: Evolution of Fourier modes during K-type transition, obtained with Simson. The solid lines are FP64 (double precision), and the black dotted lines are FP32 (single precision). a) shows the two-dimensional Fourier modes $|\hat{u}(\alpha,\beta=0)|$, starting with $\alpha=0$ on the top. (b) Three-dimensional Fourier modes $|\hat{u}(\alpha=1,\beta)|$ starting with $\beta=0$ on the top. The wall-normal maximum is shown in both cases.
@@
Figure 8: Evolution in time of amplitude of 2D modes $|\hat{u}(\alpha,\beta=0)|$ for transitional case. All simulations carried out in Neko and SSDC, except the reference case in Simson, are shown in gray. Results from full FP32 and FP64 (a), State rounding (b), and rounding of the convective term (c) are shown.
Figure 9: Profiles for the cylinder at $Re_{D}=3900$ in Neko and SSDC with rounding the convective term, the state, and running the entire solver in single and double precision. The $C_{p}$ profile in the center of the wake (a) and the wake profile at four different locations in the wake $(0.58,1.06,1.54,2.02)$ (b), the blue-shaded interval is between time averages of two low-frequency modes as described by [54].
@@
Setup Avg. time ($D/U_{\infty}$) $f_{vs}$ $\phi_{s}$ $L_{r}$ $\bar{C}_{d}$ $\bar{C}_{Pb}$
Ref. Mode L [54] $\sim$1166 0.218 87.8 1.55 0.979 0.877
Ref. Mode H [54] $\sim$1166 0.214 88.25 1.26 1.043 0.98
@@ 6 Separated flow — Cylinder at $Re_{D}=3900$
This section considers the flow around an infinite circular cylinder at $Re_{D}={U_{\infty}D}/{\nu}=3900$, where $D$ is the cylinder diameter and $U_{\infty}$ the free-stream velocity. We perform LES in PadeLibs, SSDC, and Neko with approximately 512 grid points along the cylinder boundary, and a spanwise length of $2\pi D$ with $128$ grid points. There is extensive literature on this case, both experimental [55], DNS [54] and LES [56, 57, 7]. DNS results [54] show a significant spread in the form of two distinct wake configurations, which requires a significant averaging time for accurate convergence. The results are illustrated in Figure 9, which shows the velocity profiles in the wake and pressure distribution on the cylinder surface. The velocity profiles fall within the blue-shaded band defined by the two distinct solutions. In detail, Table 4 compares the current results with the high-energy mode (Mode H), and the low-energy mode (Mode L), as described by [54]. Wall quantities such as the drag coefficient and separation angle, as well as the length of the recirculation zone agree with reference data up to the mode separation issue. Overall, the differences among the setups and solvers are comparable to the spread in the reference data. As such, for the simulations that do not diverge, results suggest that the aforementioned sources of uncertainty, i.e., averaging times, are more significant than the numerical precision when it comes to this specific test case.
Isolating the impact of lower precision may become clearer with longer averaging times; however, due to the discrepancies among multiple reference data, it is unclear whether the impact of precision can be isolated. A case with a stronger consensus among the reference data and where shorter averaging times are necessary would likely be better suited to evaluate the impact of numerical precision alone.
For a more nuanced comparison, we employ the Wasserstein distance of the probability density functions of the drag coefficient $C_{D}$. The Wasserstein distance is a metric to compare the similarity between PDFs and has previously been employed for a similar purpose (impact of rounding and floating-point precision) in [6] for climate models, where the similarity of the statistical description of the system is under consideration.
By comparing the Wasserstein distance between the drag coefficient $C_{D}$ for the different cases, the impact of averaging times is further amplified. We show the distance between the probability distributions of $C_{D}$ among the different simulations in Figure 10. In this plot, the variability for the simulations that were only carried out for 100 time units is clearly evident. For Neko, it is only for the longer averaging times that FP32 and FP64 consistently have a smaller Wasserstein distance than the poorer averaged results. While the convective FP32 might appear to differ, by comparing the longer FP64 simulation with exactly the same setup but averaged for a shorter time, we also observe a large difference that aligns with the low-frequency oscillations described in [54].
@@ 7 Compressible flow around a wing section
This test case considers the flow around a NACA-0012 airfoil at $Re_{C}={U_{\infty}C}/{\nu}=50000$, $Ma={U_{\infty}}/{c_{\infty}}=0.4$ and $\alpha=$5\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, where $C$ is the airfoil chord, $U_{\infty}$ is the free-stream velocity, $c_{\infty}$ is the speed of sound and $\alpha$ is the angle of attack. The choice of this specific configuration is intended to analyze the effect of compressibility on reduced precision computations and to consider flows with separation, transition, and curved boundaries.
The compressible Navier–Stokes solver SSDC is considered in this test case. Reference DNS calculations have been presented in the works of [58, 59]. In detail, we perform a DNS with $1080$ points on the airfoil surface and $100$ points along the spanwise direction, using a C-type grid that mimics the one from [58]. In particular, at the coordinate of maximum $C_{f}$ along the airfoil chord, $\Delta x^{+}=3.6$, $\Delta y^{+}=1.0$, and $\Delta z^{+}=6.0$. The grid extends in the wake direction for $5$ chord lengths and in the front with a radius of $7.3$; the spanwise dimension is $0.2$. The total number of points is roughly $2.58\text{\times}{10}^{8}$. Given the high computational cost, only a comparison of the double-precision and single-precision computation with CPFloat is performed.
This flow regime is characterized by a laminar separation bubble with transition and turbulent reattachment, as illustrated by plotting iso-contours of the second invariant of the velocity gradient tensor in Figure 11. The flow is initialized with a preliminary two-dimensional solution. The simulation is run in FP64 for $15$ convective time units ($C/U_{\infty}$), after which the statistics are computed for an extra $12$ convective time units for both considered precisions.
Table 5 compares the averaged integral loads and the separation bubble extension between the two SSDC simulations and the reference data. Figure 12 shows the mean pressure coefficient and skin friction coefficient on the airfoil surface. Overall, single precision computations achieve results similar to those of double precision, and both compare well to the reference. Minor differences with respect to the reference are visible in the skin friction coefficient plot, close to the leading edge on the pressure side. These can likely be attributed to a slightly different discretization of the laminar boundary layer, which is very thin in this region. Figure 13 shows mean velocity profiles and the separation bubble along the suction side for both precisions. Both computations achieve very similar results even for these quantities. The time dependence of the separation is assessed considering the probability density function of the skin friction coefficient along the suction side of the airfoil, Figure 14. In particular, the same procedure as in [58], Fig. 13c, has been applied, for example for binning. Even for this sensitive quantity, both solutions demonstrate remarkably similar results.
@@
Setup $t_{avg}(C/U_{\infty})$ $C_{L}$ $C_{D}$ $C_{D_{P}}$ $C_{D_{sf}}$ $\text{x}_{1}|_{C_{f}=0}$ $\text{x}_{2}|_{C_{f}=0}$
@@
Table 5: Scalar aerodynamic results associated with the airfoil at $Re_{C}=50000$ and $M=0.5$. Columns correspond to each setup name, associated time-averaging duration (in units of $C/U_{\infty}$), the lift coefficient $C_{L}$, the drag coefficient $C_{D}$ (split into pressure and skin friction components), and the start and end points of the separation bubble, $\text{x}_{1,2}|_{C_{f}=0}$.
@@
Figure 14: Contour plot of the PDF of the friction coefficient $C_{f}$ along the airfoil chord. The PDF is scaled, given a non-uniform bin size along $x$; more details can be found in [58] (Eq. 3.1, Fig. 13).
@@ 8 Practical experiences
This section summarizes our experiences in running our solvers natively in FP32. We comment on both performance gains (particularly on GPUs) and a range of issues that needed to be addressed for the solvers to run problem-free. While the identified problems differ among the solvers, collectively they serve as useful pointers to what should be treated with extra care when implementing an FP32-capable solver. Tools such as the Herbie project [36] could be extremely valuable in catching some of these issues during code development and making the implementation more robust to reduced precisions.
@@ 8.1 Performance improvements with FP32 over FP64
One of the most important practical outcomes from the simulation campaign was evaluating the performance impact of using FP32 instead of FP64 for the different test cases. Recall that FP32 was the only precision for which the rounding was available directly on the hardware and not emulated, thus for the other precisions such measurements were not possible to obtain.
We gained most experience from our simulations with Neko where a performance boost around $2\times$ was expected as most kernels operate in the memory-bound domain. However, it was observed that the performance improvement changed depending on whether server-grade or consumer GPUs were used, as illustrated in Table 6.
In general, it was found that a larger problem size was required to achieve a $2\times$ performance improvement on server grade GPUs when transitioning from FP64 to FP32. For the cylinder case, for example, running on the AMD Instinct MI250X GPUs, the performance was approximately $2\times$ faster with single precision. However, when a smaller turbulent channel flow case was considered, this was no longer as evident as the problem size was insufficient to hide the latency of kernel launches and oversubscribe the available computational resources.
A comparison between the server-grade Nvidia A100 and the consumer-level Nvidia RTX 4080 revealed that, for single precision simulations, the A100 achieved a performance increase of approximately $1.5\times$ over its double precision counterpart in the smaller channel flow case. In contrast, the RTX 4080 demonstrated more than double the performance in single precision, and its FP32 performance matched that of the much more expensive A100, which operates in FP64. This suggests that it is feasible to perform DNS on consumer-grade GPUs using FP32, achieving performance comparable to that of traditional FP64 simulations on high-end server GPUs. In this case, simulations executed in FP32 on the RTX 4080 produced results equivalent to those previously obtained with FP64 on the A100—within the same runtime. It is worth noting that despite the higher PF32 performance ($\pi_{\rm FP32}$) of RTX4080, A100 is still around $2\times$ faster in FP32 due to its higher memory bandwidth.
@@
RTX4080 0.72 TB/s 0.76 TFlop/s 48.7 TFlop/s $2-2.5\times$
A40 0.7 0.58 37.4 $2-2.5\times$
A100 1.56 9.7 19.4 $1.5-2\times$
@@
The second code which was used in native FP32 compilation mode was Simson. Note that only FP32 (and FP64) runs could be performed using hardware implementation (i.e. compiled for the specific precision). We nearly got the expected two-fold increase in performance, as a combination of the faster calculations and the reduced data transfer. There was an approximately 5-10% fraction of the total execution that was not affected, including IO waiting times. The time per step with Simson, compiled and run in single precision was consistently reduced to 55% on various CPU clusters, problem sizes and number of ranks. All other precisions were emulated in software which clearly increased the runtime.